1. the angle between the tangents drawn from the point (1,4)to the parabola y=4x is


2. In what ratio is the line y-x+2=0 divides the line joining the point (3,-1) & (8,9)

Grade:

1 Answers

jitender lakhanpal
12 years ago

Dear  rajan,

eqution of pair of tangent   SS1 = T2

 12(Y2 - 4X)  =(4Y-2X-2)2

WE GET   4X2 +4Y2 -16XY - 40X-16Y+4 = 0

TANß = 2√H2-AB/A+B       ABOVE EQUTION H = -8 ,A = 4,B=4

AND PUTTING THE VALUES WE GET TANß = √3  ß = 60 AND 300

QUES 2     

   LET Xm , Ym  BE THE POINTS THAT LIES ON THE LINE and divides the line joining (3, -1) (8,9) in ratio m,n

 

Xm = (3m+8n)/(m+n)     Ym = (-m+9n)/(m+n)  AS THE POINTS LIE ON THE LINE SO 

Ym-Xm+2=0

  (3m+8n)/(m+n) -(-m+9n)/(m+n) = -2

taking n common 

(3(m/n) + 8) /((m/n)+1) -(-(m/n) + 9) /((m/n)+1) = -2 

take m/n as p 

solving we get 

  p = 1/6  is the answer 

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar: Click here to download the toolbar..  

jitender 

askiitians expert